OpenStudy (anonymous):

two less than four times a number is equal to two more than three times the same number

OpenStudy (anonymous):

\[4x-2=3x+2\]

OpenStudy (anonymous):

if the number is called x, then "2 less than four times the number" is \[4x-2\] and also "two more than 3 times the number" is \[3x+2\]

OpenStudy (anonymous):

is means equal, so set them equal. solve via \[4x-2=3x+2\] \[4x-3x=2+2\]\[x=4\]

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

yw